# On the Alleged Locality in the Schrödinger Picture

### v1

@@ Abstract
Vedral claims [1] that the Schrödinger picture can describe quantum systems as locally as the Heisenberg picture, relying on a product notation for the density matrix. This reply refutes that claim. I show that the so-called ‘local factors’ in the product notation do not correspond to individual systems and therefore fail to satisfy Einsteinian locality. Furthermore, the product notation does not track where local gates are applied. Finally, I expose internal inconsistencies in the argument: if, as is also stated, Schrödinger-picture locality ultimately depends on explicit bookkeeping of all operations, then the explanatory power of the product notation is de facto undermined.
@@ 1 Motivation
In the Heisenberg picture of quantum computation [2, 3, 4], quantum systems can be fully described through local descriptions of their parts, even when the parts are entangled. A bipartite system $AB$ admits descriptors $\boldsymbol{q}_{A}$ and $\boldsymbol{q}_{B}$ that are
@@
In this reply, I demonstrate that this claim does not hold. Even in the product notation, neither the density operator nor its factors simultaneously satisfy (a) and (b), because the so-called ‘local factors’ are not local.
The central focus Ref. [1] concerns a related but more specific issue: identifying the location of phase shifts and other local gates. As is well-known, the usual Schrödinger state does not generically permit this identification. For instance, applying a phase shift (a $Z$ gate) to the first qubit of a pair in the state $\rvert\Phi^{+}\rangle=\left(\rvert 00\rangle+\rvert 11\rangle\right)/{\sqrt{2}}$  produces $({\rvert 00\rangle-\rvert 11\rangle})/{\sqrt{2}}$, which is the same state that would be obtained if the phase shift were applied to the second qubit.
@@
I will show in §3 that this, too, is false. The ‘locality in the Schrödginger picture’ suggested in the title is, therefore, misconceived. All quoted passages are from Ref. [1].
@@ 2 The Argument
First, I present
the calculation in this paper that shows that the Schrödinger picture is as Einstein local as the Heisenberg one.
@@
$\rho(0)=\frac{1}{2}(I+X_{1})(I+X_{2})\,,$
where the following notation is being used:
$X_{1}$ $=$ $X\otimes\mathds{1}\qquad Z_{1}=Z\otimes\mathds{1}$
@@
Between time 00 and $1$, a ctrl-$Z$ gate is applied, so $U(1,0)=\left(I+Z_{1}+Z_{2}-Z_{1}Z_{2}\right)/2$. This yields
$\rho(1)=\frac{1}{2}(I+X_{1}Z_{2})(I+Z_{1}X_{2})\,.$ (1)
@@
It is this form that tricks us into believing that something non-local is occurring in quantum physics. Performing the phase operation on the first qubit here leads to the state $\rho(2a)=\frac{1}{4}(I-X\otimes Z+Z\otimes X-Y\otimes Y)$, and this form just does not tell us which of the two qubits was affected (since [we] could have obtained the same state by a suitable phase kick on the second qubit). In the product notation, on the other hand, the state would become
$\rho(2a)=\frac{1}{2}(I-X_{1}Z_{2})(I+Z_{1}X_{2})\,,$
@@
Although not explicitly discussed in the article, one can verify the effect of applying a phase to the second qubit, i.e., if $U(2b,1)=Z_{2}$. The density operator would become
$\rho(2b)=\frac{1}{2}(I+X_{1}Z_{2})(I-Z_{1}X_{2})\,,$
@@
As I will explain in the next section, this is false.
@@ 3 Why the Product Notation Fails
The example is fine-tuned. I present two modifications of the scenario, both making explicit the fact that the factors in Eq. (1) do not pertain to any system. Thus, the so-called ‘local factors’ are not local in any meaningful sense.
@@
$|\Phi^{+}\rangle\langle\Phi^{+}|=\frac{1}{2}(1+X_{1}X_{2})(1+Z_{1}Z_{2})\,.$
@@ 4 Internal Contradiction
If retaining local phase information requires explicitly tracking all operations, then what is the contribution of the product notation? After all, there is also a way of using a mathematical constant, say $\pi$, to retain the local knowledge of the phase: this is done by leaving $\pi$ aside and independently keeping track of all operations executed on the system.
In short, insofar as the article asserts the locality of the factors in the product notation, it is incorrect. And where it raises the failure of the product notation, it is correct, but self-contradictory.
@@ Acknowledgements
I am grateful for insightful discussions with Samuel Kuypers, Simone Rijavec, William Schober, Vlatko Vedral and Maria Violaris, which contributed to the development of this work. These discussions took place during my visit to Wolfson College, University of Oxford, which was made possible by Vlatko Vedral’s recommendation and academic support. I especially thank him for facilitating my visit, for valuable exchanges and for including me in the scientific discourse of his research group. This work was supported by the Fonds de recherche du Québec – Nature et technologie, the Swiss National Science Foundation, and the Hasler Foundation.

### v2

@@ Abstract
Vedral claims [1] that the Schrödinger picture can describe quantum systems as locally as the Heisenberg picture, relying on a product notation for the density matrix. Here, I refute that claim. I show that the so-called ‘local factors’ in the product notation do not correspond to individual systems and therefore fail to satisfy Einsteinian locality. Furthermore, the product notation does not track where local gates are applied. Finally, I expose internal inconsistencies in the argument: if, as is also stated, Schrödinger-picture locality ultimately depends on explicit bookkeeping of all operations, then the explanatory power of the product notation is de facto undermined.
@@ 1 Motivation
In the Heisenberg picture, quantum systems can be fully described through local descriptions of their parts, even when the parts are entangled [2, 3, 4]. A bipartite system $AB$ admits descriptors $\boldsymbol{q}_{A}$ and $\boldsymbol{q}_{B}$ that are
@@
In the current reply, I demonstrate that this claim does not hold. Even in the product notation, neither the density operator nor its factors simultaneously satisfy (a) and (b), because the so-called ‘local factors’ are not local.
The central focus of Ref. [1] concerns a related but more specific issue: identifying the location of phase shifts and other local gates. As is well-known, the usual Schrödinger state does not generically permit this identification. For instance, applying a phase shift (a Pauli $Z$ gate) to the first qubit of a pair in the state $\rvert\Phi^{+}\rangle=\left(\rvert 00\rangle+\rvert 11\rangle\right)/{\sqrt{2}}$  produces $({\rvert 00\rangle-\rvert 11\rangle})/{\sqrt{2}}$, which is the same state that would be obtained if the phase shift were applied to the second qubit.
@@
I show in §3 that this, too, is false. The ‘locality in the Schrödginger picture’ suggested in the title is, therefore, misconceived. All quoted passages are from Ref. [1].
@@ 2 The Argument
First, I present the calculation of Ref. [1], which, in Vedral’s words,
shows that the Schrödinger picture is as Einstein local as the Heisenberg one.
@@
$\vspace{-4pt}\rho(0)=\frac{1}{4}(I+X_{1})(I+X_{2})\,,$
where the following notation for Pauli operators is being used:
$\vspace{-3pt}X_{1}$ $=$ $X\otimes\mathds{1}\qquad Z_{1}=Z\otimes\mathds{1}$
@@
Between time 00 and $1$, a controlled-$Z$ gate is applied, so
$U(1,0)=\operatorname{diag}(1,1,1,-1)=\left(I+Z_{1}+Z_{2}-Z_{1}Z_{2}\right)/2\,.$
This yields
$\vspace{-2pt}\rho(1)=\frac{1}{4}(I+X_{1}Z_{2})(I+Z_{1}X_{2})\,.$ (1)
@@
It is this form that tricks us into believing that something non-local is occurring in quantum physics. Performing the phase operation on the first qubit here leads to the state $\rho(2a)=\frac{1}{4}(I-X\otimes Z+Z\otimes X-Y\otimes Y)$, and this form just does not tell us which of the two qubits was affected (since [one] could have obtained the same state by a suitable phase kick on the second qubit). In the product notation, on the other hand, the state would become
$\vspace{-2pt}\rho(2a)=\frac{1}{4}(I-X_{1}Z_{2})(I+Z_{1}X_{2})\,,\vspace{-2pt}$
@@
Although not explicitly discussed in Ref. [1], one can verify the effect of applying a phase to the second qubit, i.e., if $U(2b,1)=Z_{2}$. The density operator would become
$\rho(2b)=\frac{1}{4}(I+X_{1}Z_{2})(I-Z_{1}X_{2})\,,$
@@
As I explain in the next section, this is false.
@@ 3 Why the Product Notation Fails
The example given by Verdral, and reexposed in Section 2, is fine-tuned. I present two modifications of the scenario, both making explicit the fact that the factors in Eq. (1) do not pertain to any system. Thus, the so-called ‘local factors’ are not local in any meaningful sense.
@@
$|\Phi^{+}\rangle\langle\Phi^{+}|=\frac{1}{4}(1+X_{1}X_{2})(1+Z_{1}Z_{2})\,.$
@@ 4 Internal Contradiction
If retaining local phase information requires explicitly tracking all operations, then what is the contribution of the product notation? After all, there is also a way of using a mathematical constant, for example, $\pi$, to retain the local knowledge of the phase: this is done by leaving $\pi$ aside and independently keeping track of all operations executed on the system.
In short, insofar as Ref. [1] asserts the locality of the factors in the product notation, it is incorrect. And where it raises the failure of the product notation, it is correct, but self-contradictory.
@@ Acknowledgements
I am grateful for insightful discussions with Samuel Kuypers, Simone Rijavec, William Schober, Vlatko Vedral and Maria Violaris, which contributed to the development of this work. These discussions took place during my visit to Wolfson College, University of Oxford, which was made possible by Vlatko Vedral’s recommendation and academic support. I especially thank him for facilitating my visit, for valuable exchanges and for including me in the scientific discourse of his research group.
This work was supported in part by the Fonds de recherche du Québec – Nature et technologies, the Swiss National Science Foundation, the Hasler Foundation, and the Mitacs Elevate postdoctoral fellowship in partnership with Bbox Digital.